# Evolution of reconnection flux during eruption of magnetic flux ropes

### v2

@@ Abstract
Coronal mass ejections (CMEs) are powerful drivers of space weather, with magnetic flux ropes (MFRs) widely regarded as their primary precursors. However, the variation in reconnection flux during the evolution of MFR during CME eruptions remains poorly understood. In this paper, we develop a 3D magneto-hydrodynamic model using which we explore the temporal evolution of reconnection flux during the MFR evolution using both numerical simulations and observational data. Our initial coronal configuration features an isothermal atmosphere and a potential arcade magnetic field beneath which an MFR emerges at the lower boundary. As the MFR rises, we observe significant stretching and compression of the overlying magnetic field beneath it. Magnetic reconnection begins with the gradual formation of a current sheet, eventually culminating with the impulsive expulsion of the flux rope. We analyze the temporal evolution of reconnection fluxes during two successive MFR eruptions while continuously emerging the twisted flux rope through the lower boundary. We also conduct a similar analysis using observational data from the Helioseismic and Magnetic Imager (HMI) and the Atmospheric Imaging Assembly (AIA) for an eruptive event. Comparing our MHD simulation with observational data, we find that reconnection flux play a crucial role in determination of CME speeds. From the onset to the eruption, the reconnection flux shows a strong linear correlation with the velocity. This simulation of a solar eruption provides important insights into the complex dynamics of CME initiation and progression.
1 Introduction
@@
Coronal Mass Ejections are composed of clouds of magnetized plasma that are expelled from the Sun into the heliosphere due to sudden release of free magnetic energy stored in the twisted coronal magnetic field (Chen, 2017). They are of interest due to scientific and technological reasons since CMEs can drive interplanetary shocks that energize solar particles and cause significant space weather effects in the geospace. In-situ data obtained by satellites passing through the interplanetary CMEs (ICMEs) have established that typical ICMEs have the structure of highly twisted MFR (Wang et al., 2016; Hu, 2017). Many CME models therefore incorporate a magnetic flux rope– consisting of helical field lines twisting about a central axis in the corona as the basic underlying magnetic field structure for CME precursors (Titov & Démoulin, 1999; Gibson & Fan, 2006; Duan et al., 2019; Liu, 2020; Chen, 2017). Therefore, understanding their evolution in early phase is crucial for CME studies.
@@
Although is it widely accepted that MFRs constitute the core structure of the CMEs, it remains unknown wheather MFRs exist in the solar corona before CME initiation or forms during the eruption (Chen, 2011; Patsourakos et al., 2020). Some opine that the MFR may exist before eruptions (Cheng et al., 2011), although there is no agreement on how and where an MFR might form. An alternate hypothesis is that the MFRs can emerge from beneath the photosphere (Fan, 2001; Martínez-Sykora et al., 2008; Magara, 2004; Archontis et al., 2009). Yet another proposition is that the MFRs can be built directly in the corona via shearing of magnetic footpoints and reconnection prior to the eruption (van Ballegooijen & Martens, 1989; Amari et al., 2003; Aulanier et al., 2010; Chatterjee et al., 2016). Most eruptions, particularly those originating from solar active regions (ARs), occur along magnetic polarity inversion lines (PILs) within strong field regions. Furthermore, some flare-productive ARs exhibit relatively short time interval between successive eruptions, while displaying very similar structure in CME morphology. This kind of event is referred to as homologous eruptions (Zhang & Wang, 2002). Observations indicate that the evolution of the source regions of these homologous events often display specific evolutionary pattern, including (a) continuous shearing motion (Li et al., 2010; Romano et al., 2015, 2018; Sarkar et al., 2019), (b) rotation of sunspots (Régnier & Canfield, 2006; Zhang et al., 2008), and (c) emergence of flux (Ranns et al., 2000; Nitta & Hudson, 2001; Sterling & Moore, 2001; Dun et al., 2007; Xu et al., 2017).
@@
To explain solar eruptions many theories have been suggested (Forbes et al., 2006; Shibata & Magara, 2011; MacTaggart & Hood, 2009; Schmieder et al., 2013; Aulanier, 2014; Janvier et al., 2015), but only two main models have been successful to explain homologous eruptions using magnetohydrodynamics (MHD) simulations. These are flux emergence model (Chatterjee & Fan, 2013) and the breakout model (DeVore & Antiochos, 2008). The flux emergence model suggests that new magnetic flux rises from below the solar surface, bringing with it twisted magnetic field lines created by turbulent convection that trigger eruptions (Fan, 2010). Chatterjee & Fan (2013) demonstrated repeated CMEs caused by a highly twisted flux tube emerging into the solar corona. In their simulation, a coronal flux rope partially erupted and reformed multiple times. However, the time gap between two consecutive eruptions was insufficient, the post-eruption field did not have enough time to return to a relax state before the onset of the subsequent eruption. In contrast, the breakout model features a multipolar magnetic configuration with a null point located above a central flux system which is sheared by photospheric motion. As the shearing increases, the magnetic reconnection begins at the null point. This reconnection process acts as a trigger for an eruption. When the eruption occurs, it removes the overlying magnetic field above the core that was previously restraining the core. After the eruption, the original magnetic structure is restored (Antiochos et al., 1999). Continuous shearing motions can result in repetition of such mechanisms and give rise to multiple eruptions, which are otherwise tend to be confined (DeVore & Antiochos, 2008).
Figure 1: Location of STEREO spacecraft and Earth on 2011 August 4.
@@
Although magnetic reconnection can only be indirectly observed, it is a critical process in the solar corona that results in the formation of closed magnetic loops which energizes the plasma and particles leading to sudden increase in flare radiation. The correlation suggests a connection between the CME evolution kinematics and the reconnection (Zhang et al., 2001; Temmer et al., 2008; Bein et al., 2012; Patsourakos et al., 2013). Magnetic reconnection is a ubiquitous process during coronal mass ejection (CME) events, but its precise role in the eruption sequence remains a subject of debate. It’s not definitively established whether reconnection acts as the trigger for the eruption or if it’s a result of the eruption process. Many have also utilized observations of flare emission patterns in the lower solar atmosphere as a means to indirectly study and infer the characteristics of reconnection processes occurring higher up in the corona (Forbes & Priest, 1984; Kopp & Poletto, 1986). Observing the early evolution and acceleration of coronal mass ejections (CMEs) in the lower corona presents significant challenges. As a result, many studies opt for an alternative approach. They focus on comparing two key parameters: the CME velocity, measured at several solar radii after the CME has undergone its peak acceleration, and the total magnetic flux that was reconnected during the associated flare event. Note that most of the studies have focused on at a instant after the peak phase and it is unclear weather the evolution flux is at all related to the velocity. Numerical MHD simulations have proven to be powerful tools for reproducing the time-dependent, nonlinear evolution of 3D magnetic configurations and investigating the temporal changes in flux during the early phases of magnetic flux rope (MFR) eruptions. A key property characterizing solar flares is the amount of magnetic flux passing through the reconnection sheet beneath the MFR, commonly referred to as the reconnection flux (Gopalswamy et al., 2017). The magnetic flux reconnected in the solar corona cannot be directly observed or measured. However, one can directly quantify it through its relationship with the magnetic flux encompasses by the flare ribbon or the foot points of the MFR (Kazachenko et al., 2017; Qiu et al., 2004).
Figure 2: Different phases of the filament eruption on 2011 August 04 as observed by STEREO-A. (a) Image of the Sun in STEREO-A EUVI 304 Å (̇b) COR1 image superimposed with STEREO-A EUVI 304 Å (̇c) Superimposed images of STEREO-A EUVI 304 Å , COR1 and COR2. The yellow arrows in panels (a) and (b) indicate the filament leading edge. The images are plotted using JHelioviewer (https://www.jhelioviewer.org/).
In this paper, we conduct a 3-dimensional compressible MHD simulation to study the time evolution of the reconnection flux and correlate it with the velocity of the MFR during its early evolution. We further compare our analysis with an observed event. The observational data and methods is mentioned in Section 2, the numerical setup is presented in Section 3. The results are mentioned in Section 4 and we conclude with a summary in Section 5.
2 Observation Data
We study the temporal evolution of re-connection flux and the associated CME speed during the M-class flare (SOL2011-SOL2011-08-04T03:41) which occurred in AR 11261 during 2011 August 4. During this period, STEREO-A and STEREO-B were positioned at longitudinal separations of 101 and 92 (see Figure 1), respectively, relative to the Sun-Earth line. This positioning allowed the two spacecraft to provide near-limb views of Earth-directed CMEs. On the other hand the location of AR 11261 was close to the solar disk center (between 30to 35 west) relative to Earth, offering reliable data on the solar magnetic field of the AR. This provides a unique opportunity to study the co-temporal evolution of re-connection flux, estimated from on-disk SDO data, and the speed of the associated eruption, starting from the initiation height at lower corona as observed from near-limb observations by STEREO.
Figure 4: (a) Time profile of signed cumulative reconnection flux integrated over the positive (red) and negative (blue) magnetic polarities underlying the flare ribbons. (b) Time profile of unsigned instantaneous reconnection flux integrated over the positive (red) and negative (blue) magnetic polarities underlying the flare ribbons. (c) Height time plot of the filament leading edge measured from STEREO EUVI, COR1 and COR2 for the same event as shown in Figure 2. (d) Velocity profile of the filament leading edge derived from Panel (c).
We analyze the kinematics of the eruption associated with the M-class flare on 2011 August 04, by tracking the filament leading-edge height as observed by STEREO-A. We use two instruments from the Sun Earth Connection Coronal and Heliospheric Investigation (SECCHI; Howard et al. (2008)) onboard STEREO-A to track the early evolution of the Earth-directed filament eruption, which was observed as prominence eruption from STEREO-A (see Figure 2). The instruments include the Extreme Ultraviolet Imager (EUVI; $<1.7R_{\odot}$, Wuelser et al. (2004)) and the inner coronagraph COR1 ($1.3-4R_{\odot}$, Thompson et al. (2003)).
We follow the method by (Kazachenko, 2023), to estimate the temporal profile of reconnecting flux. The observations of the flare ribbons are obtained from the Atmospheric Imaging Assembly (AIA; Lemen et al. (2012)) onboard SDO in 1600 Å channel. The associated full-disk magnetograms are obtained from the Helioseismic and Magnetic Imager (HMI; Schou et al. (2012)). During the recurrent eruptive M-class flare in AR 11261 on 2011 August 4, we identify the area swept out by the flare-ribbons observed in AIA 1600 Å images as shown in Figure 3. The half of the unsigned magnetic flux underlying the area associated with the brightening observed in AIA 1600 Å are recorded as the temporal profile of the re-connection flux. The kinematics of the filament eruption and the temporal variation in associated reconnection flux are shown in Figure 4.
3 Numerical Model
@@
In the above, $\mathbf{U}$ is the velocity field; $\mathbf{B}$ is the magnetic field; $\mathbf{A}$ is the magnetic potential; $\mathbf{J}$ is the current density; $\rho$, $p$ and $T$ are, respectively, the plasma density, pressure and temperature of the system. The viscous force is modelled as
@@
where $\nu$ is kinematic viscosity and $\mathrm{S}$ is the traceless rate-of-strain tensor. Additionally, we incorporate a semi-relativistic Boris correction into the classical MHD momentum equation to reduce numerical diffusion and avoid the need for excessively small time steps in MHD simulations, achieved by using an artificially reduced speed of light (Chatterjee, 2020).
We include radiative cooling in our model, represented in equation (4), where $\Lambda(T)$ denotes the radiative loss function as described by (Cook et al., 1989). The cooling function is plotted in Figure (5). Furthermore, we use a simplified coronal heating function that varies only with height above the solar surface. This function follows an exponential decrease as the height increases.
@@
where $\mathbf{q}_{\mathrm{sp}}=K_{sp}T^{5/2}\hat{\mathbf{b}}(\hat{\mathbf{b}}\cdot\nabla T)$, $\hat{\mathbf{b}}$ denotes the unit vector along the field direction and $K_{sp}=10^{-6}$ $\mathrm{erg\,K^{-7/2}\,cm^{-1}\,s^{-1}}$. $\tau_{sp}$ denotes a finite timescale for $\mathbf{q}_{\mathrm{cond}}$ to converge to the Spitzer heat flux and is equal to 0.1 s. Our time step varies between 0.1 - 0.3 ms.
The computational domain consists of a spherical wedge with an extent $R_{\odot}<r<6R_{\odot}$, $125\pi/12<\theta<7\pi/$ 12 and $-\pi/9.6<\phi<\pi/9.6$. It is resolved by a grid of $512\times 288\times 160$, which is non uniform in $r$ but uniform in $\theta$ and $\phi$. The grid spacing in $r$ is $dr=0.002R_{\odot}$  at the lower boundary which gradually increases in a logarithmic manner reaching $dr=0.003R_{\odot}$ at the upper boundary.
Figure 6: The twisted torus right before emerging through the bottom boundary at $r=R_{\odot}$ (in gray). The number of turns of a single field line (black) around the circular axis of the torus is shown.
Figure 7: The 3D evolution of the magnetic field of a the twisted flux rope as it emerges into the solar corona at the specified times (in hours). Red field lines represents field line anchored in the ambient arcade, while blue, green, and cyan field lines indicate field lines originating from the emerging flux region. An animation of this evolution is available in the online paper.
We consider a high-temperature coronal plasma as an ideal gas with an adiabatic constant of $\gamma=1.66$ Before the simulation starts, the domain is considered to be in hydrostatic equilibrium at a uniform temperature of $T_{0}=10^{6}$ K with a density stratification given by
$\rho=\rho_{0}\mathrm{exp}\left\{-\frac{R_{\odot}}{H_{p}}(1-\frac{R_{\odot}}{r})\right\}$ (8)
where, $\rho_{0}=1.0\times 10^{-15}$ g cm-3 is the initial density and $H_{p0}=c_{s0}/g_{r}$ denotes the initial pressure scale height. The initial atmosphere consists of a pre-existing potential field with arcade like geometry as given in Fan (2012) (see their Equation 12 and 13).
@@
which bodily transports the twisted torus radially into the domain. The major and minor radius of the torus are $0.25R_{\odot}$ and $0.042R_{\odot}$ respectively and with a field line twist rate of 0.068 rad Mm-1. The twist rate indicates the field line winds around the tube axis per unit length along the axis. The flux tube is initially located at $r_{0}=0.705{R_{\odot}}$, thus the torus is initially completely beneath the surface and then ascends bodily upward at constant speed toward the lower boundary, until it reaches a height $r_{\mathrm{stop}}$ when the emergence is stopped and a part of the torus is still inside of the computational domain. The velocity field at the lower boundary is uniformly set in the region where the emerging torus intersects tit and its zero elsewhere. The imposed emergence speed, $v_{0}$ = 2 $\mathrm{km\,s^{-1}}$, which is much smaller than the Alfven speed $v_{A}$ = 1.69 $\mathrm{Mm\,s^{-1}}$, ensuring that the emerging flux rope evolves quasi-statically during the flux emergence phase at the lower boundary. When the emergence is stopped, the velocity at the lower boundary is set to zero, with no inflows or outflows and footpoints are rigidly anchored. For the $\theta$ boundaries, a non-penetrating stress-free condition is applied for the velocity field while the magnetic field is treated with perfectly electrically conducting walls. The $\phi$ boundaries are periodic. For the top boundary, a simple outward extrapolating boundary condition is used, allowing plasma and magnetic field to flow through.
@@
In this simulation, we drive the emergence of the torus until $r_{\mathrm{stop}}=0.85R_{\odot}$ to study the reconnection flux during the evolution of the twisted flux rope in the corona which produces homologous CMEs.
4 Results

### v3

@@ Abstract
Coronal mass ejections (CMEs) are powerful drivers of space weather, with magnetic flux ropes (MFRs) widely regarded as their primary precursors. However, the variation in reconnection flux during the evolution of MFR during CME eruptions remains poorly understood. In this paper, we develop a 3D magneto-hydrodynamic model using which we explore the temporal evolution of reconnection flux during the MFR evolution using both numerical simulations and observational data. Our initial coronal configuration features an isothermal atmosphere and a potential arcade magnetic field beneath which an MFR emerges at the lower boundary. As the MFR rises, we observe significant stretching and compression of the overlying magnetic field beneath it. Magnetic reconnection begins with the gradual formation of a current sheet, eventually culminating with the impulsive expulsion of the flux rope. We analyze the temporal evolution of reconnection fluxes during two successive MFR eruptions while continuously emerging the twisted flux rope through the lower boundary. We also conduct a similar analysis using observational data from the Helioseismic and Magnetic Imager (HMI) and the Atmospheric Imaging Assembly (AIA) for an eruptive event. Comparing our MHD simulation with observational data, we find that reconnection flux play a crucial role in determination of CME kinematics. From the onset to the eruption, the rate of reconnection shows a monotonic variation with the acceleration. This simulation of a solar eruption provides important insights into the complex dynamics of CME initiation and progression.
footnotetext: $\dagger$ Now at Department of Physics & Astronomy, Georgia State University, Atlanta, GA, USA and NASA Goddard Space Flight Center, Greenbelt, MD , USAfootnotetext: $\ddagger$ Now at Department of Astronomy, Institute of Physics, Eötvös Loránd University (ELTE), Budapest, Hungary
I Introduction
@@
Coronal Mass Ejections (CMEs) are composed of clouds of magnetized plasma that are expelled from the Sun into the heliosphere due to sudden release of free magnetic energy stored in the twisted coronal magnetic field (Chen, 2017). They are of interest due to scientific and technological reasons since CMEs can drive interplanetary shocks that energize solar particles and cause significant space weather effects in the geospace. In-situ data obtained by satellites passing through the interplanetary CMEs (ICMEs) have established that typical ICMEs have the structure of highly twisted MFR (Wang et al., 2016; Hu, 2017). Many CME models therefore incorporate a magnetic flux rope– consisting of helical field lines twisting about a central axis in the corona as the basic underlying magnetic field structure for CME precursors (Titov & Démoulin, 1999; Gibson & Fan, 2006; Duan et al., 2019; Liu, 2020; Chen, 2017). Therefore, understanding their evolution in early phase is crucial for CME studies.
@@
Although is it widely accepted that MFRs constitute the core structure of the CMEs, it remains unknown wheather MFRs exist in the solar corona before CME initiation or forms during the eruption (Chen, 2011; Patsourakos et al., 2020). Some opine that the MFR could exist prior to eruptions (Cheng et al., 2011), although there is no consensus on how and where an MFR might form. An alternate hypothesis is that the MFRs can bodily emerge from the below the photosphere (Fan, 2001; Martínez-Sykora et al., 2008; Magara, 2004; Archontis et al., 2009). Yet another proposition is that the MFRs can be built directly in the corona via shearing of magnetic footpoints and reconnection prior to the eruption (van Ballegooijen & Martens, 1989; Amari et al., 2003; Aulanier et al., 2010; Chatterjee et al., 2016). MFRs can also form during the eruption (Lynch et al., 2004; Wang et al., 2017). Most eruptions, particularly those originating from solar active regions (ARs), occur along magnetic polarity inversion lines (PILs) within strong field regions. Furthermore, some flare-productive ARs exhibit relatively short time interval between successive eruptions, while displaying very similar structure in the flare emissions and CME morphology. Such kind of event has been known as homologous eruptions (Zhang & Wang, 2002). Observations indicate that the evolution of the source regions of homologous events is often characterized by continuous shearing motion (Li et al., 2010; Romano et al., 2015, 2018; Sarkar et al., 2019), sunspot rotation (Régnier & Canfield, 2006; Zhang et al., 2008), and flux emergence (Nitta & Hudson, 2001; Sterling & Moore, 2001; Ranns et al., 2000; Dun et al., 2007; Xu et al., 2017).
@@
There exists several models to explain the slow build up and abrupt release of energy in solar eruptions (Forbes et al., 2006; Shibata & Magara, 2011; Chen, 2011; Schmieder et al., 2013; Aulanier, 2014; Janvier et al., 2015). The two models widely used for homologous eruptions using magnetohydrodynamics (MHD) simulations are - the breakout reconnection models (DeVore & Antiochos, 2008) and the slow tether cutting (Fan, 2010). The breakout model invokes a multipolar magnetic configuration with a null point located above a central flux system which is sheared by photospheric motion. As the shearing increases, the magnetic reconnection begins at the null point above the newly forming flux rope. This reconnection process acts as a trigger for an eruption, When the eruption occurs, it removes the overlying magnetic field above the core that was previously restraining the core. After the eruption, the original magnetic structure is restored (Antiochos et al., 1999; Wyper et al., 2017; Lynch et al., 2008). Continuous shearing motions can result in the repetition of such mechanisms and give rise to multiple eruptions, which otherwise tend to be confined DeVore & Antiochos (2008). On the other hand, the slow tether-cutting model suggests that once a coronal flux rope is formed in the corona, its slow rise is governed by reconnection between the field lines of the twisted flux rope and the ambient field lines of the corona (Chen & Shibata, 2000; Archontis & Hood, 2008; Fan, 2010). The reconnection site usually lies below the rising flux rope. Chatterjee & Fan (2013) demonstrated repeated CMEs caused by a highly twisted flux tube emerging into the solar corona. In their simulation, a coronal flux rope partially erupted and reformed multiple times aided by partial internal reconnection between the legs of the flux rope. However, the time between eruptions was too short for the magnetic field to fully stabilize before the next eruption occurred.
@@
Although magnetic reconnection can only be indirectly observed, it is a critical process in the solar corona that forms closed loops and energizes the plasma and particles leading to impulsively enhanced flare radiation. The observed correlation between the evolution of the CME kinematics and that of the flare X-ray fluxes suggests that the CME eruption is related to the reconnection (Zhang et al., 2001; Temmer et al., 2008; Bein et al., 2012; Patsourakos et al., 2013), consistent with CME initiation models (Chen, 2011). Magnetic reconnection almost always occurs during a CME eruption yet it remains unclear whether reconnection initiates the eruption or a consequence. The morphological evolution of flare emission in the lower atmosphere has also been used to infer the reconnection process in the corona (Forbes & Priest, 1984; Kopp & Poletto, 1986). Due to the difficulty in observing CME evolution and measuring CME acceleration in the low corona, many studies instead compare the CME velocity measured at a few solar radii after the peak acceleration phase and the total magnetic flux reconnected during the flare. Note that most of the studies have focused on an instant after the peak phase (Qiu & Yurchyshyn, 2005; Pal et al., 2018; Gopalswamy et al., 2018) and it is unclear whether the reconnection flux is related to the speed during the CME evolution. Numerical MHD simulations have proven to be powerful tools for reproducing the time-dependent, nonlinear evolution of 3D magnetic configurations and investigating the temporal changes in flux during the early phases of magnetic flux rope (MFR) eruptions. A key property characterizing solar flares is the amount of magnetic flux passing through the reconnection sheet beneath the MFR, commonly referred to as the reconnection flux (Gopalswamy et al., 2017). While the reconnected flux cannot be measured directly from observations of the corona, a quantitative relationship between the reconnection flux in the corona is the magnetic flux swept by the flare ribbon or the foot points of the magnetic flux rope (Kazachenko et al., 2017; Qiu et al., 2004).
In this paper, we conduct a 3-dimensional compressible MHD simulation to study the time evolution of the reconnection flux and find the correlation with the speed of the magnetic flux rope during its early evolution. Next we aim at verifying the signatures of this correlation in observations. While, an ideal event required to estimate the reconnection flux correctly should be an on-disk event, the perfect event to calculate the ejection speed of the CME is instead a limb event. With the availability of multiple vantage points due to Stereo-A and SDO spacecraft provides an opportunity to check for particular events captured simultaneously by both spacecrafts. We analyse one of such rare events using coronagraph data from stero-A and EUV data from AIA instrument of SDO. The observational data and methods are described in § 2, the numerical setup is presented in § 3. The results are presented in § 4 and finally we summarize our work.
II Numerical Model
@@
In the above, $\mathbf{U}$ is the velocity field; $\mathbf{B}$ is the magnetic field; $\mathbf{A}$ is the magnetic potential; $\mathbf{J}$ is the current density; $\rho$, $p$ and $T$ are, respectively, the plasma density, pressure and temperature of the system; $c_{v}$ is the specific heat at constant volume. The viscous force is modelled as
@@
where $\nu$ is kinematic viscosity and $\mathrm{S}$ is the traceless rate-of-strain tensor. The coefficients $\zeta_{shock}$, represents shock viscosity which is given by
$\zeta_{shock}=\nu_{shock}\left<max_{3}[(-\nabla\cdot\mathbf{U})_{+}]\right$>
where $\left<max_{3}[(-\nabla\cdot\mathbf{U})_{+}]\right$> means that at each grid point, a value is assigned corresponding to the maximum positive flow convergence ($\nabla\cdot\mathbf{U}>0$) within three neighboring grid points along each spatial dimension. The resulting values are then smoothed using a running mean over three neighboring grid points in each coordinate direction.
Additionally, we incorporate a semi-relativistic Boris correction $F_{corr}$ into the classical MHD momentum equation to reduce numerical diffusion and avoid the need for excessively small time steps in MHD simulations, achieved by using an artificially reduced speed of light (Chatterjee, 2020).
We include radiative cooling in our model, represented in equation (4), where $\Lambda(T)$ denotes the radiative loss function as described by (Cook et al., 1989). The cooling function is plotted in Figure (1). Furthermore, we use a simplified coronal heating function that varies only with height, following an exponential decay.
@@
where $\mathbf{q}_{\mathrm{sp}}=K_{sp}T^{5/2}\hat{\mathbf{b}}(\hat{\mathbf{b}}\cdot\nabla T)$, $\hat{\mathbf{b}}$ denotes the unit vector along the field direction and $K_{sp}=10^{-6}$ $\mathrm{erg\,K^{-7/2}\,cm^{-1}\,s^{-1}}$. $\tau_{sp}$ represents a finite timescale for $\mathbf{q}_{\mathrm{cond}}$ to evolve toward the Spitzer heat flux and is set to 0.1 s. Our time step varies between 0.1 - 0.3 ms. While our setup includes most of the relevant physics, solar wind is not included at the moment.
The computational domain consists of a spherical wedge with an extent $R_{\odot}<r<6R_{\odot}$, $5\pi/12<\theta<7\pi/12$ and $\pi/9.6<\phi<\pi/9.6$. It is resolved by a grid of $512\times 288\times 160$, which is non uniform in $r$ but uniform in $\theta$ and $\phi$. The grid spacing in $r$ is $dr=0.002R_{\odot}$  at the lower boundary which gradually increases in a logarithmic manner reaching $dr=0.003R_{\odot}$ at the upper boundary. Although the lower boundary in our setup is set at $R_{\odot}$, it does not represent the photosphere; instead, it serves as the lower coronal boundary.
Figure 2: The twisted torus before emerging through the bottom boundary. The major ($r$) and minor ($a$) radius of the torus is $0.25R_{\odot}$ and $0.042R_{\odot}$. The number of turns of one of the field lines is shown in black.
Figure 3: The 3D evolution of the magnetic field of the twisted flux rope emerging into the corona at the specified times (in hours). Red field lines have footpoints in the ambient arcade, while blue (core), green (middle), and cyan (periphery) field lines originate from the emerging flux region at different distances from the flux rope axis. An animation of this evolution is available in the online article with a running time of 36 s. The period covered by the animation (in solar hours) spans from $t=19.65$ hrs to $t=40.25$ hrs from the start of the simulation. The flux rope starts emerging into the corona at $t=20.3$ hrs.
We consider a high-temperature coronal plasma as an ideal gas with an adiabatic constant of $\gamma=1.66$ . Initially, the domain is assumed to be in the state of a hydrostatic equilibrium at a uniform temperature of $T_{0}=10^{6}$ K with a density stratification given by
$\rho=\rho_{0}\mathrm{exp}\left\{-\frac{R_{\odot}}{H_{p_{0}}}(1-\frac{R_{\odot}}{r})\right\}$ (8)
Here, $\rho_{0}=1.0\times 10^{-15}$ g cm-3 is the initial density and the initial pressure scale height is given by $H_{p_{0}}=c_{s_{0}}^{2}/g_{r}$, where $c_{s_{0}}$ denotes the initial sound speed and $g_{r}$ represents the gravitational acceleration at the solar surface. The initial atmosphere consists of a pre-existing potential field with arcade like geometry as given in Fan (2012) (see their Equation 12 and 13).
@@
which bodily transports the twisted torus radially into the domain. The major and minor radius of the torus are $0.25R_{\odot}$ and $0.042R_{\odot}$ respectively (see Fig. 2) and with a field line twist rate of 0.068 rad Mm-1. The twist rate indicates the field line winds around the tube axis per unit length along the axis. The flux tube is initially located at $r_{0}=0.705{R_{\odot}}$, thus the torus is initially entirely below the surface and moves bodily upward toward the lower boundary at constant speed, until it reaches a height $r_{\mathrm{stop}}$ when the emergence is stopped and a part of the torus is still inside of the computational domain. Here $r_{0}$ denotes the position of the center of the torus at the beginning of the injection whereas $r_{\mathrm{stop}}$ represents the position of the center of the torus when the emergence stopped. The velocity field at the lower boundary is specified to be uniform in the region where the emerging torus intersects the lower boundary and is zero everywhere else. The imposed emergence speed, $v_{0}$ = 2 $\mathrm{km\,s^{-1}}$, which is much smaller than the Alfven speed $v_{A}$ = 1.69 $\mathrm{Mm\,s^{-1}}$ to ensure that the emerging flux rope is allowed to evolve quasi-statically during the flux emergence phase at the lower boundary. The method for introducing a twisted torus structure into the domain quasi-statically from the lower boundary using an electromotive force is well-known and has been used by Fan (2009, 2010, 2012, 2017) and Chatterjee & Fan (2013). It is a technique that gradually introduces the flux rope into the corona. The corona is expected to be in a force-free equilibrium state at all times. We avoid adding the flux rope suddenly into the corona as that may lead to numerical perturbations that disturb the existing equilibrium and perhaps trigger an unphysical eruption. It is desirable to discourage this possibility and to be sure that the loss of equilibrium is indeed due to MHD instabilities like torus and helical kink rather than numerical. When the emergence is stopped, the velocity at the lower boundary is set to zero, with no inflows or outflows and footpoints are rigidly anchored. For the $\theta$ boundaries, we assume a non-penetrating stress-free boundary for the velocity field and perfectly electrically conducting walls for the magnetic field. The $\phi$ boundaries are periodic. For the top boundary, we use a simple outward extrapolating boundary condition that allows plasma and magnetic field to flow through.
@@
In this simulation, we drive the emergence of the torus until $r_{\mathrm{stop}}=0.85R_{\odot}$ to study the rate of change of reconnection flux during the evolution of the twisted flux rope in the corona which produces homologous CMEs.
III Results